OpenStudy (anonymous):

Ture or False. The fraction \frac{5}{6} a terminating decimal. A. True B. False

OpenStudy (anonymous):

at the risk of repeating myself, try \(5\div 6\)

OpenStudy (anonymous):

I don't know what terminating means exactly ina decimal

OpenStudy (anonymous):

the denominator is \(6=2\times 3\) and the \(3\) will make it repeat they only terminate if the denominator is made up of 5's and 2's

OpenStudy (anonymous):

\[0.25\] terminates \[\frac{1}{3}=0.333...\] does not

OpenStudy (anonymous):

So what does terminate mean exactly?

OpenStudy (anonymous):

terminate means the non zero numbers stop, like \(0.1375\) repeating means they do not, like \(\frac{5}{6}=.83333...\)

OpenStudy (anonymous):

terminate end terminus terminator etc

OpenStudy (alexandervonhumboldt2):

if a number terminates, it means that it is ending

OpenStudy (anonymous):

ah ty